cosecant theta 9
and
Cotangent theta 3
Draw a right triangle and use the pythagoeran theorem to determine the third side and find the five remaining logarithmic functions of theta.
THANKYOU

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Logarithms have nothing whatever to do with right triangles. And what are you trying to say with "cosecant theta 9" Do you mean ?

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

cosecant theta = 9
and
Cotangent theta = 3
-----------------------------
Draw a right triangle and use the pythagoeran theorem to determine the third side and find the five remaining logarithmic functions of theta.
------
csc = r/y = 9/1
So r = 9 and y = 1
Solve for "x":
x = sqrt[r^2-y^2] = sqrt[81-1] = sqrt(80) = 4sqrt(5)
-----
sin = y/r = 1/9
cos = x/r = (4/9)sqrt(5)
tan = y/x = 1/(4sqrt(5))
----
csc = r/y = 9
sec = r/x = 9/(4sqrt(5))
cot = x/y = 4sqrt(5)
============================
cot = x/y = 3/1
So x = 3 and y = 1
r = sqrt(x^2 + y^2) = sqrt(9+1) = sqrt(10)
-----------
sin = y/r = 1/sqrt(10)
cos = x/r = 3/sqrt(10)
tan = y/x = 1/3
csc = r/y = sqrt(10)
sec = r/x = sqrt(10)/3
cot = x/y = 3/1 = 3
=========================
Cheers,
Stan H.
======================